OpenStudy (anonymous):

graph the equation by plotting 3 points -2x - 3y=-7

OpenStudy (anonymous):

Since equation is linear. You can plot in any three values for x and find the corresponding y values. Lets use x = -1, 0, and 1 -2(-1) - 3y =-7 2 - 3y = -7 y = 9/3 ------ -2(0) - 3y = -7 -3y = -7 y = 7/3 ------- -2(1) - 3y = -7 -2 - 3y = -7 -3y = -5 y = 5/3 Therefore three points = (-1, 3) & (0, 7/3) & (1, 5/3) Just a stab in the dark here hope its okay?

OpenStudy (anonymous):

it has to be whole numbers